# Efficient Individually Rational Recommender System under Stochastic Order

### v1

@@ 1. Type 1: $\textnormal{pos}(state(v))=\emptyset$, or
    Leaves of type 1 are associated with terminal states of the MDP (see Subsection 3.2). Leaves of type 2 are those whose sub-trees were pruned during the traversal. Moreover, $\textnormal{neg}(state(v))\subset\textnormal{neg}(state(parent(v)))$ holds in every such a leaf $v$, since otherwise we would have pruned its parent. Due to Observation C.1, every node $v$ with $state(v)=(Z,U_{<}\setminus\{a_{j^{*}},a_{j^{\prime}}\})$ is of type 2; therefore,
@@
is precisely the right-hand-side of Equation (19). ∎
  • Proof of Proposition B.2.Fix $Z\in\textnormal{suffix}(\sigma^{\textnormal{right}}_{\pi})$, and let $a_{j(Z)}=\operatorname*{arg\,min}_{a_{j}\in Z}\sigma^{\textnormal{right}}_{\pi}(a_{j})$. Let $T(\pi)$ denote the tree induced by $\pi$. Observe that
    Observation C.2.
@@
    Denote by $V$ the set of all nodes that satisfy the conditions of Observation C.2. Due to the way we constructed $\pi$, the paths from the root of $T(\pi)$ to any node in $V$ consist of actions that involve the arms $\{a_{i^{*}},a_{i^{\prime}},a_{j(Z)}\}\cup(U_{<}\setminus Z)$ solely; hence, we can focus on the off-path tree whose root is $s_{0}^{\prime}\stackrel{{\scriptstyle\text{def}}}{{=}}\{a_{i^{*}},a_{i^{\prime}}\}\cup(U_{<}\setminus Z)\cup\{a_{j(Z)}\}$, and the actions are precisely as in the tree induced by $\pi$ (according to the order of $\pi$). Denote this new tree by $T^{\prime}$, and let
@@
    is the probability of starting at $s_{0}^{\prime}$, and reaching a terminal node that belongs to $V^{\prime}$. ∎
Appendix D Base Cases for Lemma A.1
Proposition D.1.
@@
  • Proof of Proposition D.1.Let $\tilde{\mu}(a)\stackrel{{\scriptstyle\text{def}}}{{=}}\left|\mu(a)\right|$, and denote $U_{>}=\{a_{i_{1}}\}$ and $U_{<}=\{a_{j_{1}},\dots a_{j_{k}}\}$ for $k=\left|U_{<}\right|$. The probability of reaching the empty terminal state under any $\mathcal{P}$-valid policy is
@@
    i.e., the probability of successfully exploring $U_{<}$. Due to multiplication associativity, the above expression is invariant of the way we order its elements. Finally, by definition of $Q$, this implies that $Q(\pi,U_{>},U_{<})=Q(\rho,U_{>},U_{<})$. ∎
Proposition D.2.
@@
  • Proof of Proposition D.2.Let $\tilde{\mu}(a)\stackrel{{\scriptstyle\text{def}}}{{=}}\left|\mu(a)\right|$, and denote $U_{>}=\{a_{i_{1}},\dots a_{i_{k}}\}$ for $k=\left|U_{>}\right|$ and $U_{<}=\{a_{j_{1}}\}$. The probability of reaching the terminal state $(a_{j_{1}})$ under any $\mathcal{P}$-valid policy is
@@
  • Proof of Proposition D.3.We prove the claim by induction, with Proposition D.2 serving as the base case. Assume the claim holds for $\left|U_{<}\right|=k-1$. It is enough to show that if $\left|U_{<}\right|=k$, for any $a_{i}\in U_{>},a_{j}\in U_{<}$, $Q^{*}_{i,j}(U)=Q^{*}(U)$. Assume that $Q^{*}_{{i_{1}},{j_{1}}}(U)=Q^{*}(U)$, and fix any $a_{i^{\prime}}\in U_{>},a_{j^{\prime}}\in U_{<}$.
@@
Eq. (D) $=\boldsymbol{p}_{{i_{1}},{j_{2}}}(a_{i_{1}})\boldsymbol{p}_{{i_{2}},{j_{2}}}(a_{j_{2}})\boldsymbol{p}_{{i_{2}},{j_{1}}}(a_{j_{1}})\prod_{l=3}^{k}\boldsymbol{p}_{{i_{2}},{j_{l}}}(a_{j_{l}})$
@@
Combining this with Equation (D),
@@
This completes the proof of Proposition D.3. ∎
  • Proof of Claim 1.To ease readability, let $\tilde{\mu}_{i}\stackrel{{\scriptstyle\text{def}}}{{=}}\left|\mu(a_{i})\right|$ for every $a_{i}\in A$. It holds that
@@
  • Proof of Proposition D.4.The proof of this proposition goes along the lines of the proof of Proposition D.3, but we provide the details here for completeness. For simplicity, we let $\overline{Q}=1-Q$, and prove that for any two policies $\pi,\rho$ it holds that $\overline{Q}(\pi,U)=\overline{Q}(\rho,U)$.
    We prove the claim by induction, with Proposition D.1 serving as the base case. Assume the claim holds for $\left|U_{>}\right|=k-1$. It is enough to show that if $\left|U_{>}\right|=k$, for any $a_{i}\in U_{>},a_{j}\in U_{<}$, $Q^{*}_{i,j}(U)=Q^{*}(U)$. Assume that $Q^{*}_{{i_{1}},{j_{1}}}(U)=Q^{*}(U)$, and define a policy $\pi$ such that $\sigma^{\textnormal{right}}_{\pi}=(a_{j_{1}},a_{j_{2}})$ and $\sigma^{\textnormal{left}}_{\pi}=(a_{i_{1}},a_{i_{2}},\dots a_{i_{k}})$. Next, fix any $a_{i^{\prime}}\in U_{>},a_{j^{\prime}}\in U_{<}$.
@@
Combining Equation (D) and Claim 2, we get
@@
Notice that in the above expression, we first try to explore $a_{j_{2}}$, and only then $a_{j_{1}}$. Combining this with Equation (D),
@@
This completes the proof of Proposition D.4. ∎
  • Proof of Claim 2.To ease readability, let $\tilde{\mu}_{i}\stackrel{{\scriptstyle\text{def}}}{{=}}\left|\mu(a_{i})\right|$ for every $a_{i}\in A$. It holds that
@@
Appendix E Proof of Theorem A.1
  • Proof of Theorem A.1.Fix an arbitrary instance. We prove the claim by a two-dimensional induction on the size of $\textnormal{pos}(s),\textnormal{neg}(s)$, for states $s\in\mathcal{S}$. The base cases are
    • $\left|\textnormal{pos}(s)\right|\geq 2$ and $\left|\textnormal{neg}(s)\right|=1$ (Proposition F.1), and
    • $\left|\textnormal{pos}(s)\right|=1$ and $\left|\textnormal{neg}(s)\right|\geq 2$ (Proposition F.2),
    which we relegate to Section F. Next, assume the statement holds for all $s\in\mathcal{S}$ such that $\left|\textnormal{pos}(s)\right|\leq K_{1}$, $\left|\textnormal{neg}(s)\right|\leq K_{2}$ and $\left|\textnormal{pos}(s)\right|+\left|\textnormal{neg}(s)\right|<K_{1}+K_{2}$. Let $U\in\mathcal{S}$ denote a state with $\left|\textnormal{pos}(U)\right|=K_{1}$ and $\left|\textnormal{neg}(U)\right|=K_{2}$. For abbreviation, let $U_{>}\stackrel{{\scriptstyle\text{def}}}{{=}}\textnormal{pos}(U)=\{a_{i_{1}},a_{i_{2}},\dots,a_{i_{K_{1}}}\}$ and $U_{<}\stackrel{{\scriptstyle\text{def}}}{{=}}\textnormal{neg}(U)=\{a_{j_{1}},a_{j_{2}},\dots,a_{j_{K_{2}}}\}$, and assume the indices follow the stochastic order. Further, for every $a_{i}\in U_{>},a_{j}\in U_{<}$ let
@@
Figure 4: Illustration of the policies $\pi,\rho$ from Step 1 of Theorem A.1. Notice that the left sub-trees of $\pi$ and $\rho$ are identical.
@@
Next, we focus on the first sum of $W(\pi,U)$ in Equation (E). For every $\psi\in\Psi$, we denote the event $E^{\pi}_{\psi}$ as a shorthand for
@@
In words, $E^{\pi}_{\psi}$ is the event that the GMDP reaches the final state $(U_{<}\setminus\psi)$ or $(U_{<}\setminus(\psi\cup\{a_{j_{k}}\}))$ after starting in $s_{0}$ and following $\pi$. We use conditional expectation to simplify the summands in the first sum of Equation (E),
@@
and following the derivation in Equation (E) for $\rho$, we get
@@
Due to the fact that $I^{\pi}=I^{\rho}$ and relying on Equations (E)-(E), we have
@@
taking care of edge cases too. By renaming $\alpha^{\pi}_{\psi}$ to $\alpha^{\pi}_{\max(\psi)}$, i.e., $\alpha^{\pi}_{\max(\psi)}\stackrel{{\scriptstyle\text{def}}}{{=}}\alpha^{\pi}_{\psi}$, and rearranging Equation (E) using Equations (E) and (35),
@@
Proposition E.1.
Under Assumption 2.1,
@@
In fact, this is the only place in the proof of Theorem A.1 where we rely on Assumption 2.1. Equipped with Proposition E.1, we can make the final argument. For every $r,r\in\{1\dots,k-1\}$ let
@@
This ends the proof of Theorem A.1. ∎
Appendix F Statements for Theorem A.1
Proposition F.1.
@@
  • Proof of Proposition F.1.Denote $\textnormal{neg}(U)=\{a_{j}\}$. We can assume w.l.o.g. that the realization of all arms in $\textnormal{pos}(U)$ are non-positive, as otherwise every $\mathcal{P}$-valid policy will explore all the arms; thus, $W^{*}(U)=Q^{*}(U)\cdot\max\{0,X_{a_{j}}\}$. Finally, the Equivalence lemma suggests that $Q^{*}(U)$ is policy invariant; hence, $W(\pi,U)=W^{*}(U)$ holds for any $\mathcal{P}$-valid policy $\pi$, and in particular for $\pi=\pi^{\star}$. ∎
Proposition F.2.
@@
  • Proof of Proposition F.2.This statement is a special case of Proposition F.3 for instances satisfying Assumption 2.1. ∎
Proposition F.3.
@@
  • Proof of Proposition F.3.Denote $\textnormal{pos}(U)=U_{>}=\{a_{i}\}$ and $\textnormal{neg}(U)=U_{<}=\{a_{j_{1}},\dots a_{j_{k}}\}$ for $k=\left|U_{<}\right|$. Let $\pi$ be any right-ordered policy with the matching $\sigma^{\textnormal{right}}_{\pi}$, such that $\pi\neq\pi_{f^{*}}$. Assume that there are indices $r,l$, for $1\leq r,l\leq k$, such that $\sigma^{\textnormal{right}}_{\pi}(a_{l})<\sigma^{\textnormal{right}}_{\pi}(a_{r})$ yet $f(a_{l})<f(a_{{r}})$, for arms $a_{l},a_{r}\in U_{<}$. Moreover, if such a pair $(l,r)$ exists, assume w.l.o.g. that $\sigma^{\textnormal{right}}_{\pi}$ orders them consequentially, i.e., for every arm $a\in U$ such that $a\notin\{a_{l},a_{r}\}$, either $\sigma^{\textnormal{right}}_{\pi}(a)<\sigma^{\textnormal{right}}_{\pi}(a_{l})$ or $\sigma^{\textnormal{right}}_{\pi}(a)>\sigma^{\textnormal{right}}_{\pi}(a_{r})$.
@@
which contradicts our assumption that $f(a_{l})<f(a_{r})$. ∎
  • Proof of Claim 3.Notice that we can represent $f^{\pi}_{Z}$ as
@@
    where $a_{i(Z)}$ is the minimal element in $Z$ according to $\sigma^{\textnormal{left}}_{\pi}$. This process is similar in spirit to the proof of Proposition B.1 and is hence omitted. Then, we can mirror the same arguments for $f^{\rho}_{Z}$. Finally, the Equivalence lemma suggests that the two representations are equal. ∎
  • Proof of Proposition E.1.We address the two parts separately below.
@@
In addition, due to Assumption 2.1, $R(U_{<}\setminus(\psi\cup\{a_{j_{k}}\}))\geq R(U_{<}\setminus(\psi\cup\{a_{j_{k+1}}\}))$. Wrapping up,
@@
where the second to last inequality is due to Assumption 2.1 and independence of $(X_{a_{i}})_{i=1}^{K}$. Ultimately,
@@
This completes the proof of the second part. ∎
  • Proof of Claim 4.The proof goes along the lines of Claim 3, and is hence omitted. ∎
Appendix G Proofs of Observations and Propositions from Sections 3 and 4
  • Proof of Observation 3.1.Let $x(a_{i})>0$ for some $i\in[K]$, let $j$ be an index of unexplored arm, and let $\mathcal{I}$ be the information of the algorithm. We overload the notation $\boldsymbol{p}_{i,j}$ to acknowledge the realized value $x(a_{i})$; that is,

### v2

@@ 1. Type 1: $\textnormal{pos}(state(v))=\emptyset$, or
Leaves of type 1 are associated with terminal states of the MDP (see Subsection 3.2). Leaves of type 2 are those whose sub-trees were pruned during the traversal. Moreover, $\textnormal{neg}(state(v))\subset\textnormal{neg}(state(parent(v)))$ holds in every such a leaf $v$, since otherwise we would have pruned its parent. Due to Observation 5, every node $v$ with $state(v)=(Z,U_{<}\setminus\{a_{j^{*}},a_{j^{\prime}}\})$ is of type 2; therefore,
@@
is precisely the right-hand-side of Equation (19).
Proof (Proof of Proposition 7)
Fix $Z\in\textnormal{suffix}(\sigma^{\textnormal{right}}_{\pi})$, and let $a_{j(Z)}=\operatorname*{arg\,min}_{a_{j}\in Z}\sigma^{\textnormal{right}}_{\pi}(a_{j})$. Let $T(\pi)$ denote the tree induced by $\pi$. Observe that
Observation 6
@@
Denote by $V$ the set of all nodes that satisfy the conditions of Observation 6. Due to the way we constructed $\pi$, the paths from the root of $T(\pi)$ to any node in $V$ consist of actions that involve the arms $\{a_{i^{*}},a_{i^{\prime}},a_{j(Z)}\}\cup(U_{<}\setminus Z)$ solely; hence, we can focus on the off-path tree whose root is $s_{0}^{\prime}\stackrel{{\scriptstyle\text{def}}}{{=}}\{a_{i^{*}},a_{i^{\prime}}\}\cup(U_{<}\setminus Z)\cup\{a_{j(Z)}\}$, and the actions are precisely as in the tree induced by $\pi$ (according to the order of $\pi$). Denote this new tree by $T^{\prime}$, and let
@@
is the probability of starting at $s_{0}^{\prime}$, and reaching a terminal node that belongs to $V^{\prime}$.
Appendix 0.D Base Cases for Lemma 2
Proposition 8
@@ Appendix 0.D Base Cases for Lemma 2
Proof (Proof of Proposition 8)
Let $\tilde{\mu}(a)\stackrel{{\scriptstyle\text{def}}}{{=}}\left|\mu(a)\right|$, and denote $U_{>}=\{a_{i_{1}}\}$ and $U_{<}=\{a_{j_{1}},\dots a_{j_{k}}\}$ for $k=\left|U_{<}\right|$. The probability of reaching the empty terminal state under any $\mathcal{P}$-valid policy is
@@
i.e., the probability of successfully exploring $U_{<}$. Due to multiplication associativity, the above expression is invariant of the way we order its elements. Finally, by definition of $Q$, this implies that $Q(\pi,U_{>},U_{<})=Q(\rho,U_{>},U_{<})$.
Proposition 9
@@
Proof (Proof of Proposition 9)
Let $\tilde{\mu}(a)\stackrel{{\scriptstyle\text{def}}}{{=}}\left|\mu(a)\right|$, and denote $U_{>}=\{a_{i_{1}},\dots a_{i_{k}}\}$ for $k=\left|U_{>}\right|$ and $U_{<}=\{a_{j_{1}}\}$. The probability of reaching the terminal state $(a_{j_{1}})$ under any $\mathcal{P}$-valid policy is
@@
Proof (Proof of Proposition 10)
We prove the claim by induction, with Proposition 9 serving as the base case. Assume the claim holds for $\left|U_{<}\right|=k-1$. It is enough to show that if $\left|U_{<}\right|=k$, for any $a_{i}\in U_{>},a_{j}\in U_{<}$, $Q^{*}_{i,j}(U)=Q^{*}(U)$. Assume that $Q^{*}_{{i_{1}},{j_{1}}}(U)=Q^{*}(U)$, and fix any $a_{i^{\prime}}\in U_{>},a_{j^{\prime}}\in U_{<}$.
@@
Eq. (0.D) $=\boldsymbol{p}_{{i_{1}},{j_{2}}}(a_{i_{1}})\boldsymbol{p}_{{i_{2}},{j_{2}}}(a_{j_{2}})\boldsymbol{p}_{{i_{2}},{j_{1}}}(a_{j_{1}})\prod_{l=3}^{k}\boldsymbol{p}_{{i_{2}},{j_{l}}}(a_{j_{l}})$
@@
Combining this with Equation (0.D),
@@
This completes the proof of Proposition 10.
Proof (Proof of Claim 1)
To ease readability, let $\tilde{\mu}_{i}\stackrel{{\scriptstyle\text{def}}}{{=}}\left|\mu(a_{i})\right|$ for every $a_{i}\in A$. It holds that
@@
Proof (Proof of Proposition 11)
The proof of this proposition goes along the lines of the proof of Proposition 10, but we provide the details here for completeness. For simplicity, we let $\overline{Q}=1-Q$, and prove that for any two policies $\pi,\rho$ it holds that $\overline{Q}(\pi,U)=\overline{Q}(\rho,U)$.
We prove the claim by induction, with Proposition 8 serving as the base case. Assume the claim holds for $\left|U_{>}\right|=k-1$. It is enough to show that if $\left|U_{>}\right|=k$, for any $a_{i}\in U_{>},a_{j}\in U_{<}$, $Q^{*}_{i,j}(U)=Q^{*}(U)$. Assume that $Q^{*}_{{i_{1}},{j_{1}}}(U)=Q^{*}(U)$, and define a policy $\pi$ such that $\sigma^{\textnormal{right}}_{\pi}=(a_{j_{1}},a_{j_{2}})$ and $\sigma^{\textnormal{left}}_{\pi}=(a_{i_{1}},a_{i_{2}},\dots a_{i_{k}})$. Next, fix any $a_{i^{\prime}}\in U_{>},a_{j^{\prime}}\in U_{<}$.
@@
Combining Equation (0.D) and Claim 2, we get
@@
Notice that in the above expression, we first try to explore $a_{j_{2}}$, and only then $a_{j_{1}}$. Combining this with Equation (0.D),
@@
This completes the proof of Proposition 11.
Proof (Proof of Claim 2)
To ease readability, let $\tilde{\mu}_{i}\stackrel{{\scriptstyle\text{def}}}{{=}}\left|\mu(a_{i})\right|$ for every $a_{i}\in A$. It holds that
@@
Appendix 0.E Proof of Theorem 0.A.1
Proof (Proof of Theorem 0.A.1)
Fix an arbitrary instance. We prove the claim by a two-dimensional induction on the size of $\textnormal{pos}(s),\textnormal{neg}(s)$, for states $s\in\mathcal{S}$. The base cases are
  • $\left|\textnormal{pos}(s)\right|\geq 2$ and $\left|\textnormal{neg}(s)\right|=1$ (Proposition 13), and
  • $\left|\textnormal{pos}(s)\right|=1$ and $\left|\textnormal{neg}(s)\right|\geq 2$ (Proposition 14),
which we relegate to Section 0.F. Next, assume the statement holds for all $s\in\mathcal{S}$ such that $\left|\textnormal{pos}(s)\right|\leq K_{1}$, $\left|\textnormal{neg}(s)\right|\leq K_{2}$ and $\left|\textnormal{pos}(s)\right|+\left|\textnormal{neg}(s)\right|<K_{1}+K_{2}$. Let $U\in\mathcal{S}$ denote a state with $\left|\textnormal{pos}(U)\right|=K_{1}$ and $\left|\textnormal{neg}(U)\right|=K_{2}$. For abbreviation, let $U_{>}\stackrel{{\scriptstyle\text{def}}}{{=}}\textnormal{pos}(U)=\{a_{i_{1}},a_{i_{2}},\dots,a_{i_{K_{1}}}\}$ and $U_{<}\stackrel{{\scriptstyle\text{def}}}{{=}}\textnormal{neg}(U)=\{a_{j_{1}},a_{j_{2}},\dots,a_{j_{K_{2}}}\}$, and assume the indices follow the stochastic order. Further, for every $a_{i}\in U_{>},a_{j}\in U_{<}$ let
@@
Figure 4: Illustration of the policies $\pi,\rho$ from Step 1 of Theorem 0.A.1. Notice that the left sub-trees of $\pi$ and $\rho$ are identical.
@@
Next, we focus on the first sum of $W(\pi,U)$ in Equation (0.E). For every $\psi\in\Psi$, we denote the event $E^{\pi}_{\psi}$ as a shorthand for
@@
In words, $E^{\pi}_{\psi}$ is the event that the GMDP reaches the final state $(U_{<}\setminus\psi)$ or $(U_{<}\setminus(\psi\cup\{a_{j_{k}}\}))$ after starting in $s_{0}$ and following $\pi$. We use conditional expectation to simplify the summands in the first sum of Equation (0.E),
@@
and following the derivation in Equation (0.E) for $\rho$, we get
@@
Due to the fact that $I^{\pi}=I^{\rho}$ and relying on Equations (0.E)-(0.E), we have
@@
taking care of edge cases too. By renaming $\alpha^{\pi}_{\psi}$ to $\alpha^{\pi}_{\max(\psi)}$, i.e., $\alpha^{\pi}_{\max(\psi)}\stackrel{{\scriptstyle\text{def}}}{{=}}\alpha^{\pi}_{\psi}$, and rearranging Equation (0.E) using Equations (0.E) and (35),
@@
Proposition 12
Under Assumption 1,
@@
In fact, this is the only place in the proof of Theorem 0.A.1 where we rely on Assumption 1. Equipped with Proposition 12, we can make the final argument. For every $r,r\in\{1\dots,k-1\}$ let
@@
This ends the proof of Theorem 0.A.1.
Appendix 0.F Statements for Theorem 0.A.1
Proposition 13
@@
Proof (Proof of Proposition 13)
Denote $\textnormal{neg}(U)=\{a_{j}\}$. We can assume w.l.o.g. that the realization of all arms in $\textnormal{pos}(U)$ are non-positive, as otherwise every $\mathcal{P}$-valid policy will explore all the arms; thus, $W^{*}(U)=Q^{*}(U)\cdot\max\{0,X_{a_{j}}\}$. Finally, the Equivalence lemma suggests that $Q^{*}(U)$ is policy invariant; hence, $W(\pi,U)=W^{*}(U)$ holds for any $\mathcal{P}$-valid policy $\pi$, and in particular for $\pi=\pi^{\star}$.
Proposition 14
@@
Proof (Proof of Proposition 14)
This statement is a special case of Proposition 15 for instances satisfying Assumption 1.
Proposition 15
@@
Proof (Proof of Proposition 15)
Denote $\textnormal{pos}(U)=U_{>}=\{a_{i}\}$ and $\textnormal{neg}(U)=U_{<}=\{a_{j_{1}},\dots a_{j_{k}}\}$ for $k=\left|U_{<}\right|$. Let $\pi$ be any right-ordered policy with the matching $\sigma^{\textnormal{right}}_{\pi}$, such that $\pi\neq\pi_{f^{*}}$. Assume that there are indices $r,l$, for $1\leq r,l\leq k$, such that $\sigma^{\textnormal{right}}_{\pi}(a_{l})<\sigma^{\textnormal{right}}_{\pi}(a_{r})$ yet $f(a_{l})<f(a_{{r}})$, for arms $a_{l},a_{r}\in U_{<}$. Moreover, if such a pair $(l,r)$ exists, assume w.l.o.g. that $\sigma^{\textnormal{right}}_{\pi}$ orders them consequentially, i.e., for every arm $a\in U$ such that $a\notin\{a_{l},a_{r}\}$, either $\sigma^{\textnormal{right}}_{\pi}(a)<\sigma^{\textnormal{right}}_{\pi}(a_{l})$ or $\sigma^{\textnormal{right}}_{\pi}(a)>\sigma^{\textnormal{right}}_{\pi}(a_{r})$.
@@
which contradicts our assumption that $f(a_{l})<f(a_{r})$.
Proof (Proof of Claim 3)
Notice that we can represent $f^{\pi}_{Z}$ as
@@
where $a_{i(Z)}$ is the minimal element in $Z$ according to $\sigma^{\textnormal{left}}_{\pi}$. This process is similar in spirit to the proof of Proposition 6 and is hence omitted. Then, we can mirror the same arguments for $f^{\rho}_{Z}$. Finally, the Equivalence lemma suggests that the two representations are equal.
Proof (Proof of Proposition 12)
We address the two parts separately below.
@@
In addition, due to Assumption 1, $R(U_{<}\setminus(\psi\cup\{a_{j_{k}}\}))\geq R(U_{<}\setminus(\psi\cup\{a_{j_{k+1}}\}))$. Wrapping up,
@@
where the second to last inequality is due to Assumption 1 and independence of $(X_{a_{i}})_{i=1}^{K}$. Ultimately,
@@
This completes the proof of the second part.
Proof (Proof of Claim 4)
The proof goes along the lines of Claim 3, and is hence omitted.
Appendix 0.G Proofs of Observations and Propositions from Sections 3 and 4
Proof (Proof of Observation 1)
Let $x(a_{i})>0$ for some $i\in[K]$, let $j$ be an index of unexplored arm, and let $\mathcal{I}$ be the information of the algorithm. We overload the notation $\boldsymbol{p}_{i,j}$ to acknowledge the realized value $x(a_{i})$; that is,